OpenStudy (anonymous):

HomePride manufactures solid oak racks for displaying baseball equipment and karate belts. They usually sell six times as many baseball racks as karate-belt racks. The net profit is $3 from each baseball rack and $5 from each karate-belt rack. If the company wants a total profit of $46,000, how many of each type of rack should they sell? Write a system of two equations to solve the problem.

OpenStudy (ash2326):

@yagirlmack Have you understood the problem?

OpenStudy (anonymous):

i dont know how to answer this

OpenStudy (ash2326):

Let the no. of baseball racks sold be x, and no. of karate-belt racks sold be y. We are given the information "They usually sell six times as many baseball racks as karate-belt racks" so \[x=6y\] do you get this part ?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

"The net profit is $3 from each baseball rack and $5 from each karate-belt rack. If the company wants a total profit of $46,000" \[Total\ profit= 3\times x+5 \times y\] We are given that company wants a profit of $46,000 so \[3x+5y=46000................(2)\] We had found one equation earlier \[x=6y....................(1)\] Can you use equation (1) and (2) to find x and y ?

OpenStudy (anonymous):

you can use equation 2

OpenStudy (anonymous):

i think, idk

OpenStudy (ash2326):

\[3x+5y=46000\] We have x=6y, let's substitute for x \[3(6y)+5y=46000\] \[18y+5y=46000\] Can you find y from this?

OpenStudy (anonymous):

yes y=2000

OpenStudy (ash2326):

Good :D Now find x from this

OpenStudy (anonymous):

how?

OpenStudy (ash2326):

You have the equation \[x=6y\]

OpenStudy (anonymous):

so i would do x=6(2000)

OpenStudy (ash2326):

yes

OpenStudy (anonymous):

x=12000

OpenStudy (anonymous):

what do i do from here?

OpenStudy (ash2326):

The question is solved :D no. of baseball racks sold is x, and no. of karate-belt racks sold is y.

OpenStudy (anonymous):

okay thank you sooo much :D

OpenStudy (ash2326):

Welcome :D